OpenStudy (anonymous):

solve the following equation: y"-4y'+13y=0 y(0)=-1 y'(0)=2

OpenStudy (anonymous):

Do you know its homo characteristic equation?

OpenStudy (anonymous):

i want to learn about these type of equations,what do you mean homo characteristic, Chloro ?

OpenStudy (anonymous):

when you see the right side = 0, It's homogeneous type!

OpenStudy (anonymous):

how then is it solved ?

OpenStudy (anonymous):

Solve for its root as second degree polynomial

OpenStudy (anonymous):

is not this a differential equation of the second order ?

OpenStudy (anonymous):

second order DE => solve by second degree polynomial!

OpenStudy (anonymous):

can you show me how it is done ?

OpenStudy (anonymous):

r^2 - 4r + 13 = 0

OpenStudy (turingtest):

^this is called the "characteristic polynomial" of the DE

OpenStudy (anonymous):

@KANNYTE can you solve it?

OpenStudy (anonymous):

are'n t we going to form a differential equation first?

OpenStudy (anonymous):

The DE goal is find y !

OpenStudy (anonymous):

i don't understand ,can you please explain ?

OpenStudy (anonymous):

Can you find the root for this: r^2 - 4r + 13 = 0

OpenStudy (anonymous):

yes i can

OpenStudy (turingtest):

@KANNYTE are you really wanting to know why solving this quadratic will help us solve our problem? are you just wanting to know why? because chlorophyll has provided the correct method

OpenStudy (anonymous):

Pls show your roots here!

OpenStudy (anonymous):

yes i want to know why

OpenStudy (anonymous):

@KANNYTE follow the instruction, then: Can you show the root for this: r^2 - 4r + 13 = 0

OpenStudy (turingtest):

we make an assumption that the solution will be some linear combination of\[y=e^{rt}\]it then follows that \[y'=re^{rt}\]and\[y''=r^2e^{rt}\]so under our assumption our eqn becomes\[r^2e^{rt}-4re^{rt}+13=0\]since \(e^{rt}\) can never be zero, that means the polynomial \[r^2-4r+13=0\]hence, if we find r, we know what our solution will look like.

OpenStudy (anonymous):

r=2+6i and r=2-6i

OpenStudy (anonymous):

Should you resolve it?

OpenStudy (turingtest):

getting complex roots means we have to use Euler's formula:\[e^{it}=\cos t+i\sin t\]using some trig identities this means that when our solutions to the characteristic polynomial are a complex conjugate of the form\[\lambda\pm \mu i\]we know that our solution will be of the form\[y=c_1e^{\lambda}\cos(\mu t)+c_2e^{\lambda t}sin(\mu t)\]

OpenStudy (turingtest):

if you want to see exactly how we get that formula read this link: http://tutorial.math.lamar.edu/Classes/DE/ComplexRoots.aspx

OpenStudy (turingtest):

typo above the expected solution will have the form\[y(t)=c_1e^{\lambda t}\cos(\mu t)+c_2e^{\lambda t}\sin(\mu t)\]

OpenStudy (anonymous):

so what will be the final solution of the equation?

OpenStudy (anonymous):

Then, use initial values f(0) and f'(0( => C1 and C2

OpenStudy (turingtest):

please don't just ask for the final answer, I'm not going to let it be that easy for you. If that is what you want learn to use wolfram alpha. chlorophyll and I have given you more than enough info to solve this problem I believe.

OpenStudy (turingtest):

note: wolfram alpha is not the best when it comes to DE's fair warning

OpenStudy (anonymous):

@TuringTest Thanks for fulfill the asker's inquirement =)

OpenStudy (turingtest):

likewise :)

OpenStudy (anonymous):

i need your help so that i can solve this, i plead you to be patient with me guys,what you have been telling me earlier is like greek to me,can you explain from the basics of differential equations ? pliz

OpenStudy (turingtest):

do you know what an integrating factor is? can you solve a linear (first order) DE ?

OpenStudy (anonymous):

yes i can

OpenStudy (turingtest):

I think your roots are wrong; they should be\[r=2\pm3i=\lambda\pm\mu i\]

OpenStudy (turingtest):

I told you that characterisitic polynomials of this form have a general solution\[y(t)=c_1e^{\lambda t}\cos(\mu t)+c_1e^{\lambda t}\sin(\mu t)\]plug in the values for \(\lambda\) and \(\mu\) that you found after that, find plug in t=0 to find y(0) and y'(0), which will give you a system of equation which can be solved for c1 and c2, just as in the linear cases

OpenStudy (turingtest):

I have to go, I hope I helped a bit check out that link I gave you for further explanation good luck!